OpenStudy (anonymous):

Show that among all rectangles with a given perimeter, the square has the largest area?

OpenStudy (zarkon):

let L and W be the length and with of a rectangle. we have 2W+2L=p (p some fixed number) area \[A=L\cdot W\] solve for W in fitst equation \[W=\frac{p-2L}{2}\] then \[A=L\frac{p-2L}{2}\] \[A=\frac{1}{2}(Lp-2L^2)\] \[\frac{dA}{dL}=\frac{1}{2}(p-4L)\] set equal to zero ..solve \[\frac{1}{2}(p-4L)=0\Rightarrow L=\frac{p}{4}\] then \[W=\frac{p-2L}{2}=\frac{p-2\frac{p}{4}}{2}=\frac{p-p/2}{2}=\frac{p/2}{2}=p/4\] thus W=L and our rectangle is really a square.

OpenStudy (zarkon):

one should also check that this is truly a max (could use 1st or 2nd derivative test...do this for all the problems you needed to do)

OpenStudy (anonymous):

sorry, didn't catch that?

OpenStudy (zarkon):

we need to determine if we found a maximum or minimum. this process will find both. the 1st or 2nd D-test will tell us if we found a max or min

OpenStudy (anonymous):

ohh okay okay, got it! ;) thank you sooo much, you're a lifesaver! ;)

OpenStudy (zarkon):

using the 2nd D-test \[\frac{d^2A}{dL^2}=\frac{1}{2}(-4)=-2\] this is true for all values of L even our critical number p/4 since the 2nd D is negative at our critical value we have found a max by the 2nd D-test

OpenStudy (zarkon):

np

OpenStudy (dumbcow):

penpal this looks good. a square always maximizes area given a perimeter

OpenStudy (dumbcow):

notice how its kinda the same process

OpenStudy (anonymous):

yeah, i noticed, thanks for looking into it. you rock!